# Towards Zero-Shot Text-To-Speech for Arabic Dialects

### v1

@@ Abstract
Khai Duy Doan duy.doan@mbzuai.ac.ae                        Abdul Waheed abdul.waheed@mbzuai.ac.ae                        Muhammad Abdul-Mageed muhammad.mageed@ubc.ca
Mohamed bin Zayed University of Artificial Intelligence / Abu Dhabi, UAE
@@ 3 XTTS Model
XTTS Eren and The Coqui TTS (2021)  [^6] We use XTTSv2 and refer to it as XTTS throughout our study. is a multilingual multispeaker speech synthesis model that supports Arabic (ar) along with English (en), and multiple other languages.
@@ 4.2 Methods
  • Fine-tuned without dialect token: we fine-tune the model with the same parameters, vocabulary as the pre-trained model.
  • Fine-tuned with dialect token: we add a list of 22 dialect tokens (21 dialects and MSA) to the vocabulary so the vocabulary size increases from 6681 to 6703 entries. To provide the model with the dialect information, we insert the dialect token after the [lang] token: s1, s2, …, sk, [bots], [lang], [dialect] t1, t2, …, tn, [eots], [boas], a1, a2, …, a_l, [eoas]. In the embedding layer, we initialize the weights for new tokens by sampling from the standard normal distribution.
@@ 4.3 Evaluation
Speaker Similarity. We use speaker embedding cosine similarity (SECS) as the quantitative metric evaluation. To compute the speaker embedding vector of speech audio, we employ the Resemblyzer package  [^12] https://github.com/resemble-ai/Resemblyzer. Afterwards, we get the cosine similarity score between two speaker embedding vectors, which is in the range [-1, 1], where a larger value indicates a higher similarity of input vectors. For each sample synthesis, we randomly pick a reference utterance, which has at least 5-second duration length, from the same speaker. We run the experiment three times and report the average score.
@@ 6 Limitations
7 Conlusion

### v3

@@ Abstract
Khai Duy Doanξ       Abdul Waheedξ       Muhammad Abdul-Mageedξ,γ,λ ξMBZUAI      γThe University of British Columbia      λ Invertible AI {duy.doan,abdul.waheed}@mbzuai.ac.ae        muhammad.mageed@ubc.ca
@@ 3 XTTS Model
XTTS Eren and The Coqui TTS (2021)  [^6] We refer to XTTSv2 simply as XTTS in our study. is a multilingual multispeaker speech synthesis model that supports Arabic (ar) along with English (en), and multiple other languages.
@@ 4.2 Methods
Fine-tuned without dialect token. We fine-tune the model with the same parameters, vocabulary as the pre-trained model.
Fine-tuned with dialect token. We add a list of 22 dialect tokens (21 dialects and MSA) to the vocabulary so the vocabulary size increases from 6681 to 6703 entries. To provide the model with the dialect information, we insert the dialect token after the [lang] token: s1, s2, …, sk, [bots], [lang], [dialect] t1, t2, …, tn, [eots], [boas], a1, a2, …, a_l, [eoas]. In the embedding layer, we initialize the weights for new tokens by sampling from the standard normal distribution.
@@ 4.3 Evaluation
Speaker Similarity. We use speaker embedding cosine similarity (SECS) as the quantitative metric evaluation. To compute the speaker embedding vector of speech audio, we employ the Resemblyzer package. Afterwards, we get the cosine similarity score between two speaker embedding vectors, which is in the range [-1, 1], where a larger value indicates a higher similarity of input vectors. For each sample synthesis, we randomly pick a reference utterance, which has at least 5-second duration length, from the same speaker. We run the experiment three times and report the average score.
@@ 7 Limitations
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^13] https://alliancecan.ca and UBC Advanced Research Computing-Sockeye. [^14] https://arc.ubc.ca/ubc-arc-sockeye